The boy in question 1 now throws his ball horizontally out of a high window with a velocity of 3.1 m s-1. How long will it take to reach the ground 18m below?



1
Expert's answer
2021-04-26T16:48:31-0400

"v_x=3 .1~\frac ms,~v_y=0~\frac ms,"

"h=v_yt+\frac{gt^2}{2}=\frac{gt^2}{2}\implies"

"t=\sqrt{\frac{2h}{g}}=1.9~s."